# On variants of Chebyshev’s conjecture

### v2

@@ 1.1. Results on the Riemann zeta function
Let $\pi(x)$ be the number of prime numbers $\leq x$. According to the prime number theorem, the leading term in the asymptotic behavior of $\pi(x)$ as $x\to\infty$ is the logarithmic integral ${\rm li}(x)=\int_{0}^{x}dt/\log t$. The difference $\pi(x)-{\rm li}(x)$ had long been conjectured to be negative for all $x$. This conjecture was disproven in 1914 by Littlewood, who showed that it changes sign infinitely often [8]. However, Rubinstein and Sarnak [19] proved that the logarithmic density of $x$ for which $\pi(x)-{\rm li}(x)$ is negative is close to one assuming the Riemann Hypothesis (RH) holds – that is, all nontrivial zeros of $\zeta(s)$ have real part $1/2$ – and the set of positive imaginary parts of the nontrivial zeros of the Riemann zeta function $\zeta(s)$ is linearly independent over $\mathbb{Q}$. Furthermore, as mentioned in Pintz [16] and as proved by Johnston [10], the RH is equivalent to the negativity
$\int_{2}^{x}(\pi(t)-{\rm li}(t))\,dt<0\quad(x>2).$ (1.1)
@@
$\psi_{1/2}(x):=\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}$ (1.3)
@@
$\psi_{1/2}(x)-2\sqrt{x}=-\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+o(1)$ (1.4)
in (2.6) below, together with Littlewood’s result [8, Theorem 5.6]
@@
we can show that $\psi_{1/2}(x)-2\sqrt{x}$ changes sign infinitely often as well as $\pi(x)-{\rm li}(x)$, where the sum $\sum_{\rho}$ runs over all nontrivial zeros of $\zeta(s)$ counted with multiplicity, and is defined as $\sum_{\rho}=\lim_{T\to\infty}\sum_{|\Im(\rho)|\leq T}$. The Landau symbol $f(x)=o(g(x))$ means that $|f(x)|/g(x)$ tends to zero as $x\to\infty$. Furthermore, $f(x)=\Omega_{+}(g(x))$ (resp. $f(x)=\Omega_{-}(g(x))$) means that $\limsup_{x\to\infty}f(x)/g(x)>0$ (resp. $\liminf_{x\to\infty}f(x)/g(x)<0$), and $f(x)=\Omega_{\pm}(g(x))$ means that both $f(x)=\Omega_{+}(g(x))$ and $f(x)=\Omega_{-}(g(x))$ hold. By the second functional equation in (2.4), if $\rho$ is a nontrivial zero of $\zeta(s)$, then so is $\bar{\rho}$. Therefore, both $\sum_{\rho}x^{\rho-1/2}/\rho$ and $\sum_{\rho}x^{\rho-1/2}/(\rho-1/2)$ are real valued.
@@
holds, $(\zeta^{\prime}/\zeta)(1/2)=2.68609\dots$ (see (3.1) and (3.2) below), and the sum on the right-hand side is bounded under the RH by $\sum_{\rho}|\rho|^{-2}<\infty$ (see §2.2). This bias in the difference $\psi_{1/2}(x)-2\sqrt{x}$ is also a sufficient condition for the RH as in the case of $\pi(x)-{\rm li}(x)$ and $\vartheta(x)-x$:
@@
    $\lim_{x\to\infty}\left[\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\log x}\right]=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right).$ (1.6)
The prime on the sum is unnecessary on the left-hand side of (1.5), since $\log(x/n)$ vanishes when $n=x$. Taking the average
@@
into account, the inequality in (1.5) can be rewritten as an inequality for the sum of $\Lambda(n)$ with a weight that is approximately one on average:
@@
The equivalence between the first and second statements in Theorem 1 is quite similar to that between the RH and either (1.1) or (1.2). However, an analog of (1.6) for $\pi(x)-\mathrm{li}(x)$ or $\vartheta(x)-x$ does not appear to have been previously noted in the literature. Moreover, it should be mentioned that Theorem 1 was not discovered as an analog of (1.1); rather, it arose in connection with the screw functions discussed below.
@@
$g_{0}(t)$ $=\sum_{n\leq\exp(|t|)}\frac{\Lambda(n)}{\sqrt{n}}(|t|-\log n)-4(e^{t/2}+e^{-t/2}-2),$ (1.7)
@@
$\xi(s):=\frac{1}{2}\,s(s-1)\pi^{-s/2}\Gamma(s/2)\zeta(s).$ (1.8)
@@
$g_{0}(\log x)=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\left(\sqrt{x}+\frac{1}{\sqrt{x}}-2\right),$ (1.9)
@@
Let $g_{0}(t)$ be the non-archimedean part of $g_{\zeta}(t)$ defined in (1.7). Then, the RH holds if and only if there exists $t_{0}>0$ such that $g_{0}(t)\geq 0$ holds for all $t\geq t_{0}$.
Noting (1.4), if we replace $4\sqrt{x}$ in (1.5) with $2\sum_{n\leq x}\Lambda(n)/\sqrt{n}$ and express $2$ as $\log\exp(2)$, we obtain the following as an analog of Theorem 1.
@@
$\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}~\leq~0$ (1.10)
@@
$\lim_{x\to\infty}\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)$ (1.11)
@@
The right-hand side of (1.11) is negative (see (3.2) below). Therefore, (1.11) implies that there exists $x_{0}\geq 2$ such that (1.10) holds for all $x\geq x_{0}$,
For the sum on the left-hand side of (1.12), the best known estimate under the RH is
@@
[9, pp. 82–85], but it is not enough to prove (1.10) for large $x$, where the Vinogradov symbol $f(x)\ll g(x)$ mean that there exists a positive constant $M$ such that $|f(x)|\leq Mg(x)$ holds for a prescribed range of $x$. We use the the Landau symbol $f(x)=O(g(x))$ in the same meaning. In the above sense, (1.10) and (1.11) may provide a condition stronger than the RH. For comparison, we note that Montgomery’s conjecture
@@
[14, p.16] implies the RH and the estimate (1.12). The equivalence of (1.11) and (1.12) can be viewed as analogous to Akatsuka [2, Theorem 1], which in turn parallels Conrad [5, Theorem 1.3] for $\zeta(s)$. Taking the logarithm of [2, Corollary 4.4] and using [2, Lemma 2.1], Akatsuka’s result asserts that
@@
is equivalent to the validity of the RH and (1.12), where $\zeta(1/2)=1.46035\cdots$. Recently, Akatsuka showed that the equation obtained by replacing the upper limit $\sqrt{x}$ of the integral on the left-hand side with $\sqrt{\vartheta(x)}$ is equivalent to the RH ([3, Theorem 3]).
@@
$\lim_{x\to\infty}\sum_{n\leq x}a_{n}\left(1-\frac{\log n}{\log x}\right)=a$
(cf. [15, p.158]). From (3.1) below, $\{\Lambda(n)/\sqrt{n}\}$ cannot be Riesz summable of order one even if the RH is assumed. However, Theorem 2 asserts that if we suppose a condition stronger than the RH, then a slightly modified Riesz sum of order one has a limit, and conversely, the existence of such a limit implies the RH.
If we rewrite (1.10) as
$\sum_{n\leq x}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{x}{n}\quad\leq~\sum_{x<n\leq xe^{2}}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{n}{x}$ (1.13)
@@ 1.2. Results related to Chebyshev’s conjecture
For any pair of integers $q$ and $a$ with $(q,a)=1$, let $\pi(x;q,a)$ denote the number of primes $p\leq x$ such that $p\equiv a$ mod $q$. The leading term in the asymptotic behavior of $\pi(x,q,a)$ as $x\to\infty$ is ${\rm li}(x)/\varphi(q)$, where $\varphi(q)$ is the Euler totient function. As in the case of the difference $\pi(x)-{\rm li}(x)$, Chebyshev’s conjecture that $\pi(x;4,1)-\pi(x;4,3)$ is always nonpositive was disproved in [8], but it has been shown that the logarithmic density of the set of $x>2$ for which $\pi(x;4,1)-\pi(x;4,3)$ is negative is close to one [19].
@@
for integers $q$ and $a$ with $(q,a)=1$. Then, a negative bias is observed in the difference
@@
where $\chi_{4}$ is the non-principal Dirichlet character modulo four. Furthermore, we find that $\psi_{1/2}(x;4,1)-\psi_{1/2}(x;4,3)$ is negative on average if and only if all zeros of the Dirichlet $L$-function $L(s,\chi_{4})$ in the critical strip $0\leq\Re(s)\leq 1$ lie on the line $\Re(s)=1/2$. The latter is nothing but the Generalized Riemann Hypothesis (GRH) for $L(s,\chi_{4})$ holds.
@@
    $\lim_{x\to\infty}\sum_{n\leq x}\frac{\Lambda(n)\chi_{4}(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi_{4}\right).$ (1.15)
Unlike [5, Theorem 1.3], [2, Theorem 1], and Theorem 2, which include claims stronger than the GRH, it should be noted that (1.15) is equivalent to the GRH for $L(s,\chi_{4})$. On the other hand, similar to Theorem 1, the nonpositivity in (1.14) can be interpreted as a property of the non-archimedean part of the screw function associated with $L(s,\chi_{4})$ studied in [22], but we omit the details, as they may be easily inferred.
@@
$\lim_{x\to\infty}\sum_{p>2}(-1)^{(p-1)/2}\cdot e^{-p/x}=-\infty.$ (1.16)
The sum on the left-hand side of (1.16) can be written as
@@
Here $e^{-p/x}\to 1$ as $x\to\infty$ termwise. Therefore, Chebyshev’s conjecture (1.16) suggests that there are “more” primes of the form $p=4n+3$ than those of the form $p=4n+1$. Subsequently, Hardy and Littlewood [8] and Landau [13] proved that (1.16) and
$\lim_{x\to\infty}\sum_{p>2}(-1)^{(p-1)/2}\log p\cdot e^{-p/x}=$-$\infty$ (1.17)
are equivalent to the GRH for $L(s,\chi_{4})$. Knapowski and Turán [11] showed that replacing the weight function $\exp(-p/x)$ in (1.16) and (1.17) with $\exp(-(\log(p/x))^{2})$ is likewise equivalent to the GRH for $L(s,\chi_{4})$. Fujii [7] proved that when $\exp(-p/x)$ is replaced with $\exp(-(p/x)^{\alpha})$ (resp. $2K_{0}(2\sqrt{p/x})$), the result in (1.16) (resp. (1.17)) is equivalent to the GRH for $L(s,\chi_{4})$ if $0<\alpha<4.19$, and he conjectured that this would hold for general $\alpha>0$ as well. Platt and Trudgian [18] extends Fujii’s result for (1.16) to $0<\alpha<20.4$. Aoki and Koyama [1] studies the case where $\exp(-p/x)$ is replaced by $p^{-s}$ ($s\geq 0$) in (1.16). In addition, a few other weight functions were considered in the works of Hardy–Littlewood [8] and Fujii [7]. Based on the above developments and Theorem 3, considering the weight function $\sqrt{x/p}\log(x/p)$ in (1.17) leads to the following result.
@@
$\lim_{x\to\infty}\sum_{2<p\leq x}(-1)^{(p-1)/2}\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\infty.$ (1.18)
@@
$\sum_{2<p\leq x}(-1)^{(p-1)/2}\sqrt{\frac{x}{p}}~\sim~\frac{1}{2}\sqrt{x}\log\log x\quad\text{as $x\to\infty$}$ (1.19)
holds if and only if the GRH for $L(s,\chi_{4})$ holds and (1.12) is satisfied for the nontrivial zeros of $L(s,\chi_{4})$ by combining [1, Theorem 1.1, (1.3)] and [5, Theorem 6.2]. According to the proof of Theorem 4, the GRH for $L(s,\chi_{4})$ is equivalent to the left-hand side of (1.18) being asymptotically equal to $(\sqrt{x}/4)(\log x)^{2}$; thus, unlike (1.19), it does not yield a result stronger than the GRH. However, an analog of Theorem 4 for $\zeta(s)$ yields a conclusion stronger than the RH, similar to Theorem 2.
@@
$\lim_{x\to\infty}\sum_{p\leq xe^{2}}\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\infty.$ (1.20)
Conversely, if the RH holds and (1.12) is valid as $x\to\infty$, then (1.20) holds.

### v3

@@ 1.1. Results on the Riemann zeta function
Let $\pi(x)$ be the number of prime numbers $\leq x$. According to the prime number theorem, the leading term in the asymptotic behavior of $\pi(x)$ as $x\to\infty$ is the logarithmic integral ${\rm li}(x)=\int_{0}^{x}dt/\log t$. The difference $\pi(x)-{\rm li}(x)$ had long been conjectured to be negative for all $x>2$. This conjecture was disproven in 1914 by Littlewood, who showed that it changes sign infinitely often [8]. However, Rubinstein and Sarnak [19] proved that the logarithmic density of $x$ for which $\pi(x)-{\rm li}(x)$ is negative is close to one assuming the Riemann Hypothesis (RH) holds – that is, all nontrivial zeros of $\zeta(s)$ have real part $1/2$ – and the set of positive imaginary parts of the nontrivial zeros of the Riemann zeta function $\zeta(s)$ is linearly independent over $\mathbb{Q}$. Furthermore, as mentioned in Pintz [16] and as proved by Johnston [10], the RH is equivalent to the negativity
$\int_{2}^{x}(\pi(t)-{\rm li}(t))\,dt<0\quad(x>2).$ (1)
@@
$\psi_{1/2}(x):=\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}$ (3)
@@
$\psi_{1/2}(x)-2\sqrt{x}=-\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+o(1)$ (4)
in (36) below, together with Littlewood’s result [8, Theorem 5.6]
@@
we can show that $\psi_{1/2}(x)-2\sqrt{x}$ changes sign infinitely often as well as $\pi(x)-{\rm li}(x)$, where the sum $\sum_{\rho}$ runs over all nontrivial zeros of $\zeta(s)$ counted with multiplicity, and is defined as $\sum_{\rho}=\lim_{T\to\infty}\sum_{|\Im(\rho)|\leq T}$. The Landau symbol $f(x)=o(g(x))$ means that $|f(x)|/g(x)$ tends to zero as $x\to\infty$. Furthermore, $f(x)=\Omega_{+}(g(x))$ (resp. $f(x)=\Omega_{-}(g(x))$) means that $\limsup_{x\to\infty}f(x)/g(x)>0$ (resp. $\liminf_{x\to\infty}f(x)/g(x)<0$), and $f(x)=\Omega_{\pm}(g(x))$ means that both $f(x)=\Omega_{+}(g(x))$ and $f(x)=\Omega_{-}(g(x))$ hold. By the second functional equation in (34), if $\rho$ is a nontrivial zero of $\zeta(s)$, then so is $\bar{\rho}$. Therefore, both $\sum_{\rho}x^{\rho}/\rho$ and $\sum_{\rho}x^{\rho-1/2}/(\rho-1/2)$ are real valued.
@@
holds, $(\zeta^{\prime}/\zeta)(1/2)=2.68609\dots$ (see (46) and (47) below), and the sum on the right-hand side is bounded under the RH by $\sum_{\rho}|\rho|^{-2}<\infty$ (see §2.2). This bias in the difference $\psi_{1/2}(x)-2\sqrt{x}$ is also a sufficient condition for the RH as in the case of $\pi(x)-{\rm li}(x)$ and $\vartheta(x)-x$:
@@
    $\lim_{x\to\infty}\left[\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)-\frac{4\sqrt{x}}{\log x}\right]=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right).$ (6)
The prime on the sum is unnecessary on the left-hand side of (5), since $\log(x/n)$ vanishes when $n=x$. Taking the average
@@
into account, the inequality in (5) can be rewritten as an inequality for the sum of $\Lambda(n)$ with a weight that is approximately one on average:
@@
The equivalence between the first and second statements in Theorem 1 is quite similar to that between the RH and either (1) or (2). However, an analog of (6) for $\pi(x)-\mathrm{li}(x)$ or $\vartheta(x)-x$ does not appear to have been previously noted in the literature. Moreover, it should be mentioned that Theorem 1 was not discovered as an analog of (1); rather, it arose in connection with the screw functions discussed below.
@@
$g_{0}(t)$ $=\sum_{n\leq\exp(|t|)}\frac{\Lambda(n)}{\sqrt{n}}(|t|-\log n)-4(e^{t/2}+e^{-t/2}-2),$ (7)
@@
$\xi(s):=\frac{1}{2}\,s(s-1)\pi^{-s/2}\Gamma(s/2)\zeta(s).$ (8)
@@
$g_{0}(\log x)=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\left(\sqrt{x}+\frac{1}{\sqrt{x}}-2\right),$ (9)
@@
Let $g_{0}(t)$ be the non-archimedean part of $g_{\zeta}(t)$ defined in (7). Then, the RH holds if and only if there exists $t_{0}>0$ such that $g_{0}(t)\geq 0$ holds for all $t\geq t_{0}$.
Noting (4), if we replace $4\sqrt{x}$ in (5) with $2\sum_{n\leq x}\Lambda(n)/\sqrt{n}$ and express $2$ as $\log\exp(2)$, we obtain the following as an analog of Theorem 1.
@@
$\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}~\leq~0$ (10)
@@
$\lim_{x\to\infty}\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)$ (11)
@@
The right-hand side of (11) is negative (see (47) below). Therefore, (11) implies that there exists $x_{0}\geq 2$ such that (10) holds for all $x\geq x_{0}$,
For the sum on the left-hand side of (12), the best known estimate under the RH is
@@
[9, pp. 82–85], but it is not enough to prove (10) for large $x$, where the Vinogradov symbol $f(x)\ll g(x)$ mean that there exists a positive constant $M$ such that $|f(x)|\leq Mg(x)$ holds for a prescribed range of $x$. We use the the Landau symbol $f(x)=O(g(x))$ in the same meaning. In the above sense, (10) and (11) may provide a condition stronger than the RH. For comparison, we note that Montgomery’s conjecture
@@
[14, p.16] implies the RH and the estimate (12). The equivalence of (11) and (12) can be viewed as analogous to Akatsuka [2, Theorem 1], which in turn parallels Conrad [5, Theorem 1.3] for $\zeta(s)$. Taking the logarithm of [2, Corollary 4.4] and using [2, Lemma 2.1], Akatsuka’s result asserts that
@@
is equivalent to the validity of the RH and (12), where $\zeta(1/2)=1.46035\cdots$. Recently, Akatsuka showed that the equation obtained by replacing the upper limit $\sqrt{x}$ of the integral on the left-hand side with $\sqrt{\vartheta(x)}$ is equivalent to the RH ([3, Theorem 3]).
@@
$\lim_{x\to\infty}\sum_{n\leq x}a_{n}\left(1-\frac{\log n}{\log x}\right)^{k}=a$
(cf. [15, p.158]). From (46) below, $\{\Lambda(n)/\sqrt{n}\}$ cannot be Riesz summable of order one even if the RH is assumed. However, Theorem 2 asserts that if we suppose a condition stronger than the RH, then a slightly modified Riesz sum of order one has a limit, and conversely, the existence of such a limit implies the RH.
If we rewrite (10) as
$\sum_{n\leq x}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{x}{n}\quad\leq~\sum_{x<n\leq xe^{2}}\Lambda(n)\sqrt{\frac{x}{n}}\log\frac{n}{x}$ (13)
@@ 1.2. Results related to Chebyshev’s conjecture
For any pair of integers $q$ and $a$ with $(q,a)=1$, let $\pi(x;q,a)$ denote the number of primes $p\leq x$ such that $p\equiv a$ mod $q$. The leading term in the asymptotic behavior of $\pi(x;q,a)$ as $x\to\infty$ is ${\rm li}(x)/\varphi(q)$, where $\varphi(q)$ is the Euler totient function. As in the case of the difference $\pi(x)-{\rm li}(x)$, Chebyshev’s conjecture that $\pi(x;4,1)-\pi(x;4,3)$ is always nonpositive was disproved in [8], but it has been shown that the logarithmic density of the set of $x>2$ for which $\pi(x;4,1)-\pi(x;4,3)$ is negative is close to one [19].
@@
for integers $q$ and $a$ with $(q,a)=1$. This is a Chebyshev function version of the weighted counting function $\pi_{1/2}(x;q,a)$ studied in Aoki and Koyama [1]. Then, a negative bias is observed in the difference
@@
where $\chi_{4}$ is the non-principal Dirichlet character modulo four. Furthermore, we find that $\psi_{1/2}(x;4,1)-\psi_{1/2}(x;4,3)$ is negative on average if and only if all zeros of the Dirichlet $L$-function $L(s,\chi_{4})$ in the critical strip $0\leq\Re(s)\leq 1$ lie on the line $\Re(s)=1/2$. The latter is nothing but the Generalized Riemann Hypothesis (GRH) for $L(s,\chi_{4})$.
@@
    $\lim_{x\to\infty}\sum_{n\leq x}\frac{\Lambda(n)\chi_{4}(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi_{4}\right).$ (15)
Unlike [5, Theorem 1.3], [2, Theorem 1], and Theorem 2, which include claims stronger than the GRH, it should be noted that (15) is equivalent to the GRH for $L(s,\chi_{4})$. On the other hand, similar to Theorem 1, the nonpositivity in (14) can be interpreted as a property of the non-archimedean part of the screw function associated with $L(s,\chi_{4})$ studied in [22], but we omit the details, as they may be easily inferred.
@@
$\lim_{x\to\infty}\sum_{p>2}(-1)^{(p-1)/2}\cdot e^{-p/x}=-\infty.$ (16)
The sum on the left-hand side of (16) can be written as
@@
Here $e^{-p/x}\to 1$ as $x\to\infty$ termwise. Therefore, Chebyshev’s conjecture (16) suggests that there are “more” primes of the form $p=4n+3$ than those of the form $p=4n+1$. Subsequently, Hardy and Littlewood [8] and Landau [13] proved that (16) and
$\lim_{x\to\infty}\sum_{p>2}(-1)^{(p-1)/2}\log p\cdot e^{-p/x}=$-$\infty$ (17)
are equivalent to the GRH for $L(s,\chi_{4})$. Knapowski and Turán [11] showed that replacing the weight function $\exp(-p/x)$ in (16) and (17) with $\exp(-(\log(p/x))^{2})$ is likewise equivalent to the GRH for $L(s,\chi_{4})$. Fujii [7] proved that when $\exp(-p/x)$ is replaced with $\exp(-(p/x)^{\alpha})$ (resp. $2K_{0}(2\sqrt{p/x})$), the result in (16) (resp. (17)) is equivalent to the GRH for $L(s,\chi_{4})$ if $0<\alpha<4.19$, and he conjectured that this would hold for general $\alpha>0$ as well. Platt and Trudgian [18] extend Fujii’s result for (16) to $0<\alpha<20.4$. In addition, a few other weight functions were considered in the works of Hardy–Littlewood [8] and Fujii [7]. Based on the above developments and Theorem 3, considering the weight function $\sqrt{x/p}\log(x/p)\mathbf{1}_{p\leq x}$ in (17) leads to the following result.
@@
$\lim_{x\to\infty}\sum_{2<p\leq x}(-1)^{(p-1)/2}\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\infty.$ (18)
@@
$\sum_{2<p\leq x}(-1)^{(p-1)/2}\sqrt{\frac{x}{p}}~\sim~\frac{1}{2}\sqrt{x}\log\log x\quad\text{as $x\to\infty$}$ (19)
holds if and only if the GRH for $L(s,\chi_{4})$ holds and (12) is satisfied for the nontrivial zeros of $L(s,\chi_{4})$ by combining [1, Theorem 1.1, (1.3)] and [5, Theorem 6.2]. According to the proof of Theorem 4, the GRH for $L(s,\chi_{4})$ is equivalent to the left-hand side of (18) being asymptotically equal to $(\sqrt{x}/4)(\log x)^{2}$; thus, unlike (19), it does not yield a result stronger than the GRH. However, an analog of Theorem 4 for $\zeta(s)$ yields a conclusion stronger than the RH, similar to Theorem 2.
@@
$\lim_{x\to\infty}\sum_{p\leq xe^{2}}\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\infty.$ (20)
Conversely, if the RH holds and (12) is valid as $x\to\infty$, then (20) holds.